OpenStudy (grace111):

MEDAL + FAN FOR BEST ANSWER !! which one has the larger value??? a) 6m or 150cm b) 900p or £ 10 c)2.5km or 2000m d) 1 1/4 hours or 65minutes

OpenStudy (grace111):

Its for tomz plzzz help :)

OpenStudy (amoodarya):

6m=6*100 cm=600 cm vs 150cm so 600 cm is larger 2.5 km =2.5 *1000 m=2500 m vs 2000 m 11/4 hours=60min +1/4(60min)=60+15 min=75 min vs 65 min